OpenStudy (anonymous):

Factor completely. 20x^6+81x^3+81

OpenStudy (anonymous):

Is there another way to do without ac method

OpenStudy (cwrw238):

not sure what you mean by the ac method

Nnesha (nnesha):

AC ----> any calculator :)

OpenStudy (anonymous):

It's quadratic?

OpenStudy (cwrw238):

yea you could change it to a quadratic by letting y = x^3 then factor 20y^2 + 81y + 81

OpenStudy (cwrw238):

factor this then substitute back x^3 for y

OpenStudy (anonymous):

I'm aware of the quadratic part where I treat it like a quadratic but my question is how can I find factors of 1620 without a calculator. It's such a large number

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

And I wasn't given a calcular and I had very little time

Nnesha (nnesha):

start dividing 1620 by 20 or 21

Nnesha (nnesha):

when you divide 1620/20 you get 81 but 81 and 20 is 101 so you can have idea that you need to divide by some bigger number

Nnesha (nnesha):

then try 1620/30 you get 54 54 + 30 equal 84 so close to81 then i will jump to 36 1620/36 = 45 and 36 + 45 = 81 looks right

Nnesha (nnesha):

well this is how i will this problem if i get it in exam :) :(((

Nnesha (nnesha):

solve ***

OpenStudy (anonymous):

Yeh. THIS would take some time if I hD only 25 mins for 20 questions

Nnesha (nnesha):

sound like SAT question :P

OpenStudy (cwrw238):

another way is to look at the factors of 20 and 81 and see how you can get 2 numbers to add up to 81